### v1

# What are they talking about? Benchmarking Large Language Models for Knowledge-Grounded Discussion Summarization

Abstract
In this work, we investigate the performance of LLMs on a new task that requires combining discussion with background knowledge for summarization. This aims to address the limitation of outside observer confusion in existing dialogue summarization systems due to their reliance solely on discussion information. To achieve this, we model the task output as background and opinion summaries and define two standardized summarization patterns. To support assessment, we introduce the first benchmark comprising high-quality samples consistently annotated by human experts and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. We evaluate LLMs under structured-prompt and self-reflection paradigms. Our findings reveal: (1) LLMs struggle with background summary retrieval, generation, and opinion summary integration. (2) Even top LLMs achieve less than average performance across both patterns. (3) Current LLMs lack adequate self-evaluation and self-correction capabilities for this task.111Our benchmark is available at zhouweixiao/KGDS
What are they talking about?
Benchmarking Large Language Models for
Knowledge-Grounded Discussion Summarization
Weixiao Zhou Junnan Zhu Gengyao Li Xianfu Cheng Xinnian Liang Feifei Zhai Zhoujun Li††thanks: Corresponding Author State Key Laboratory of Complex & Critical Software Environment, Beihang University State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, CAS University of Chinese Academy of Sciences ByteDance Fanyu AI Laboratory wxzhou@buaa.edu.cn junnan.zhu@nlpr.ia.ac.cn
1 Introduction
Dialogue summarization aims to distill main topics and interactions from a dialogue into a concise and faithful summary Jia et al. (2023); Rennard et al. (2023); Kirstein et al. (2024). Current methods typically rely exclusively on dialogue content across dataset construction Gliwa et al. (2019); Chen et al. (2021); Zhu et al. (2021); Hu et al. (2023), summarization methodologies Lin et al. (2022); Zhou et al. (2023); Tian et al. (2024), and evaluation paradigms Wang et al. (2022); Gao and Wan (2022); Zhu et al. (2023); Gao et al. (2023); Tang et al. (2023, 2024b); Ramprasad et al. (2024). These efforts implicitly assume that "Information in the dialogue is sufficiently complete to support generation of a clearly understandable summary for outside observers."
However, we find this assumption has fundamental limitations and often breaks down, particularly in discussions involving shared background knowledge. Such discussions exhibit two main traits: (1) Information Omission and Implicit Reference: Participants naturally skip mutually known details and frequently use pronouns or phrases to refer to entities or facts within the contextual background. (2) Personal Opinion: Unlike straightforward information interactions in general dialogues, these discussions focus on exchanging viewpoints, with participants expressing personal opinions from various positions or perspectives (see Figure 1). These characteristics make understanding heavily reliant on background knowledge. Consequently, traditional dialogue summarization pattern maintains this dependency, leading to confusion for outsiders unfamiliar with the context, leaving them wondering: "What are they talking about?".
To address this issue, we introduce Knowledge-Grounded Discussion Summarization (KGDS), a novel task specifically designed to combine shared background knowledge with discussion content to produce observer-friendly summaries. We determine that an effective summary must achieve two complementary objectives: (1) bridge the knowledge gap between outsiders and participants by providing necessary background-supporting information; (2) present clear personal opinions of the participants with clarified implicit references. We formalize these requirements by modeling the task output as two components: background summary and opinion summary. The background summary, which retrieves or condenses relevant contextual information, can be either extractive or abstractive, while the opinion summary integrates clarified viewpoints and is inherently abstractive. Thus, we define two standardized summarization patterns: EBS-AOS and ABS-AOS (refer to Figure 1).
To systematically investigate the challenges of KGDS and evaluate LLM capabilities in this task, we make two significant contributions. First, we construct the first KGDS benchmark containing high-quality, multi-domain samples from realistic news discussions through strict consistency-controlled annotation protocols. Each sample comprises structured shared background knowledge, multi-turn human discussion, and expert-annotated evaluation components at multiple granularities. Second, we propose a novel hierarchical evaluation framework with fine-grained and interpretable metrics (as shown in Figure 1). Specifically: (1) For extractive background summary, we utilize retrieval metrics to evaluate recall, precision, and f1-score of supporting paragraphs. (2) For abstractive background summary, inspired by the minimal granularity of atomic facts Liu et al. (2023b) and the excellent human-alignment of LLM-based fact verification Wei et al. (2024), we assess quality via verification of key supporting fact inclusion and nonsupporting fact exclusion, proposing three metrics that multi-dimensionally quantify background coverage, focus, and summary overall performance. (3) For abstractive opinion summary, we evaluate quality by verifying coverage of clear atomic opinions, where each opinion is minimized and implicit references clarified. For uncovered opinions, we define five error types and perform fine-grained error detection. (4) For pattern overall performance, we use geometric mean for evaluation, which aligns with the KGDS workflow and reflects the joint fulfillment rate of both sub-summaries.
We evaluate leading LLMs under single-turn structured-prompt Li et al. (2023) and multi-turn self-reflection Shinn et al. (2023) paradigms. Our analysis reveals several critical insights: (1) LLMs struggle with precision-recall trade-off in background summary retrieval; (2) LLMs frequently miss key facts and retain irrelevant facts in background summary generation; (3) LLMs encounter difficulties with clarifying implicit references in opinion summary integration; (4) Even advanced LLMs achieve less than average performance across both patterns; (5) LLMs lack sufficient self-evaluation and -correction capabilities for KGDS.
Our contributions are as follows: (1) We introduce a novel KGDS task and establish two standardized summarization patterns. (2) We construct the first benchmark for this task and propose a comprehensive hierarchical evaluation framework with interpretable metrics. (3) Our extensive evaluation reveals specific challenges that current LLMs face with KGDS, providing valuable guidance for future improvements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
2 Related Work
Dialogue Summarization.
Dialogue summarization is widely used in practical applications Lin et al. (2021); Hu et al. (2023). Earlier approaches concentrated on feature modeling Chen and Yang (2021); Fang et al. (2022) and pre-training methods Zou et al. (2021); Zhong et al. (2022a). The emergence of LLMs has greatly enhanced dialogue summarization Wang et al. (2023a); Tian et al. (2024); Lu et al. (2025); Zhu et al. (2025), with recent research focusing on evaluating their effectiveness in this field Liu et al. (2024b); Ramprasad et al. (2024); Tang et al. (2024b). However, these studies often overlook the importance of background knowledge.
Knowledge-Grounded Dialogue Generation.
The task involves generating responses that interact with external knowledge sources like documents Wang et al. (2023b, 2024), databases Zhou et al. (2018); Zhang et al. (2020), or retrieval-augmented systems Lewis et al. (2020); Chen et al. (2024). The main distinctions between this task and our KGDS are: (1) Conversation characteristics: The former focuses on knowledge-intensive dialogues, while KGDS deals with background-sparse discussions. (2) Task objectives: The former enhances response quality through knowledge, whereas KGDS utilizes background for complementary summarization.
Summarization Evaluation.
Earlier similarity-based metrics like Rouge Lin (2004), BERTScore Zhang et al. (2019), and MoverScore Zhao et al. (2019) are coarse-grained and often misaligned with human judgment. NLI- and QA-based approaches improve faithfulness evaluation by checking factual consistency Goyal and Durrett (2020); Laban et al. (2022); Zha et al. (2023) or generating relevant questions Wang et al. (2020); Fabbri et al. (2022); Zhong et al. (2022b), but they are limited to faithfulness and require specialized training. Recently, LLM-based reference-free evaluators Liu et al. (2023a); Chen et al. (2023); Wang et al. (2023c); Shen et al. (2023); Fu et al. (2024) have been used, yet they still lack interpretability and fine-grained assessment. To address these, some studies Song et al. (2024a, 2025); Lee et al. (2024); Yang et al. (2024); Scirè et al. (2024); Wan et al. (2024) utilize atomic fact units Liu et al. (2023b); Min et al. (2023) and LLM-based claim verification Tang et al. (2024a); Wei et al. (2024); Song et al. (2024b), achieving more fine-grained evaluation. Among these, FineSurE Song et al. (2024a) is most relevant to our work. However, its metrics lack dimensional consistency (Completeness is measured at the fact-level, while Conciseness is sentence-level), limiting them to isolated sub-aspect evaluation. In contrast, our coverage and focus metrics are both atomic-fact-level, allowing direct F1-score computation for overall quality assessment while maintaining interpretability. Our method can theoretically be adapted for general query-focused summarization, requiring only the maintenance of inner and outer fact sets.
Summarization Benchmark.
Current research often evaluates LLMs across various summarization tasks, including dialogue Tang et al. (2024b), news Goyal et al. (2022); Zhang et al. (2024), multi-document Huang et al. (2024), multilingual Ye et al. (2024), and query-focused Yang et al. (2023); Liu et al. (2024a). Moreover, some studies investigate inherent issues such as hallucinations Belem et al. (2024); Bao et al. (2024) and biases Ravaut et al. (2024); Chhabra et al. (2024). Unlike these, our work uniquely focuses on LLMs within KGDS, evaluating their capabilities in background retrieval, generation, and opinion integration.
3 Task Formulation
EBS-AOS Pattern.
The output comprises an extractive background summary (EBS) and an abstractive opinion summary (AOS):
where is the shared background knowledge of participants, is the discussion grounded in , and is an instruction for the EBS-AOS pattern. is the LLM-driven summarization system. is the EBS, defined as extractive background supporting chunks for from , where chunks represent text sequences with a predefined granularity (e.g., sentences, paragraphs, sections, or window-sliding segments). is the AOS, defined as clear personal opinions of the participants with clarified implicit references.
ABS-AOS Pattern.
The output contains an abstractive background summary (ABS) and an abstractive opinion summary (AOS):
where , , , and retain their definitions from Eq. (1). The instruction is specific to the ABS-AOS pattern. is the ABS, defined as abstractive background supporting information for from .
4 Benchmark
4.1 Preliminary
Scenario Setting.
We establish the benchmark scenario as a two-participant discussion about news content for two reasons: (1) news discussions are highly prevalent in daily communication, and (2) news summarization is a representative subfield of automatic summarization research.
News Collection.
We collect multi-domain (i.e., business, sports, and world) event-rich news articles from Google News222https://news.google.com as shared background knowledge. To ensure data credibility and avoid potential contamination in LLMs Deng et al. (2023); Li and Flanigan (2024), we control the time span of news between Mar and Sep 2024. Additionally, we retain the original paragraph structure and define paragraph-as-chunk as the minimum granularity under the EBS-AOS pattern (see Table 4 in Appendix B for more statistics).
Expert Recruitment.
We recruit four PhD candidates specializing in NLP to engage in our annotation work. Each pair of experts form a collaborative group to conduct the full-process annotation.
4.2 Discussion Construction
The construction follows the sequence of read, understand, then discuss. Initially, for each news article, we require two experts to independently read and thoroughly understand its content. The purpose of this preliminary step is to ensure knowledge consistency. Subsequently, we ask them to engage in a discussion to exchange viewpoints. This process is open-ended, meaning the initiator of the discussion is chosen at random, and the discussion topics can encompass any events, facts, or sub-information contained within the news article (see Table 5 in Appendix B for discussion statistics).
4.3 Annotation and Evaluation for EBS
Back-Supporting Paragraphs Annotation.
For each paragraph in the news article, we ask the two experts involved in the discussion: "Is this paragraph the background-supporting paragraph for discussion?" and require them to answer with either "yes" or "no". We then apply strict consistency control, retaining only paragraphs consistently annotated as "yes" (i.e., back-supporting) or "no" (i.e., back-nonsupporting). For inconsistent annotations (i.e., one "yes" and one "no"), we determine them to be boundary paragraphs and remove from the original news article (see Appendix B for annotation statistics). Our investigation in Section §6.1 shows that LLMs, with an average win rate of , prefer expert-consistently-annotated EBS over their own extractions, highlighting the excellent quality of our annotations.
Evaluation Metrics.
EBS assessment involves matching LLM-extracted paragraph indices. We use information retrieval metrics: Back-Supporting Paragraph (BSP) Recall, Precision, and F1-score (refer to Appendix C for metric details).
4.4 Annotation and Evaluation for ABS
Key Back-Supporting Atomic Facts Annotation.
This process consists of two steps: atomic fact decomposition and expert annotation. We use GPT-4o333Version: gpt-4o-2024-08-06 to decompose back-supporting paragraphs into candidate atomic facts, following the principles of indivisibility, independence, and declarativity (the instruction is provided in Appendix D.2). Next, we ask two experts for each fact: "Is this fact the key background-supporting fact for discussion?" and require them to answer either "yes" or "no". We employ the same consistency control to determine key back-supporting atomic facts (i.e., two "yes") and non-key facts (i.e., one or zero "yes" responses) (refer to Appendix B for annotation statistics).
Back-Nonsupporting Atomic Facts Annotation.
This process is automated and includes two steps: atomic fact decomposition and conflicting fact masking. We employ the same approach to decompose back-nonsupporting paragraphs into initial atomic facts. Next, we mask444The masking process is consistent with fact verification, and the conflicting facts are not considered in ABS evaluation. facts that can be inferred from back-supporting paragraphs to obtain back-nonsupporting atomic facts. Conflicts typically occur when identical or similar facts appear in both back-supporting and -nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking such facts, we construct non-overlapping back-supporting and -nonsupporting fact sets, avoiding any external verification problems (see Appendix B for annotation statistical data).
Fine-Grained Multi-Dimensional Metrics.
Let denote the LLM-generated ABS, the set of key back-supporting atomic facts, the set of back-nonsupporting atomic facts. The LLM-based fact verification function is defined as:
where and represent whether the fact can be inferred from .
Metric . KBSAF Recall
Quantifies ABS coverage of key back-supporting atomic facts:
Metric . KBSAF Precision
Measures ABS focus on key back-supporting atomic facts:
Metric . KBSAF F1-score
Evaluates ABS overall performance by comprehensively considering both coverage and focus:
4.5 Annotation and Evaluation for AOS
Clear Atomic Opinions Annotation.
The annotations are created entirely by experts and involve three steps: main opinion extraction, implicit reference clarification, and clear opinion atomization. First, we require experts to extract the main opinions from their respective utterances in the discussion. Next, we ask them to identify referential pronouns and phrases within these opinions and clarify them (i.e., through anaphora resolution or information supplementation) using referents (e.g., entities, facts, or events) from the news article to produce clear opinions. Finally, we instruct the experts to decompose these clear opinions into atomic opinion units, adhering to the principles of indivisibility and independence (see Appendix B for detailed annotation statistics).
Fine-Grained Quantification Metric.
Let be the LLM-generated AOS, and be the set of clear atomic opinions. The LLM-based opinion verification function is defined as:
where and indicate whether the opinion can be inferred from .
Metric . CAO Recall
Assesses AOS coverage of clear atomic opinions:
Fine-Grained Error Detection.
Unlike query-focused ABS, the integrativity of AOS determines that it is almost impossible to exhaust the space of erroneous atomic opinions. As a result, defining precision and F1-score for AOS evaluation is impractical. However, we can identify why certain opinions are not covered by AOS through fine-grained error detection. Specifically, for opinions that cannot be inferred from AOS, we define five error types and conduct LLM-based error classification: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion (detailed error definitions are provided in Appendix D.7).
4.6 Overall Performance Evaluation
We use the Geometric Mean to assess pattern-level overall performance for two reasons: (1) Its multiplicative property reflects the joint fulfillment rate of the background and opinion summaries, emphasizing the holistic synergy between these two components. (2) The root normalization ensures the dimensional consistency of the component metrics. Specifically, we define the overall performance metrics for the EBS-AOS and ABS-AOS patterns as:
| Model Name | KGDS BenchMark (single-turn structured-prompt and multi-turn self-reflection) | |||||
| EBS-AOS Pattern | ABS-AOS Pattern | |||||
| BSP (R-P-F1) | CAO (R) | OP (GM) | KBSAF (R-P-F1) | CAO (R) | OP (GM) | |
| GPT-4o | 73.39↑1.56-88.10↑0.34-78.12↑1.27 | 76.18↑0.54 | 76.24↑1.04 | 63.57↓1.08-61.73↑1.62-58.34↑0.60 | 69.42↓0.21 | 61.09↑0.47 |
| GPT-4-turbo | 73.03↓7.72-87.42↑1.67-77.14↓3.80 | 72.73↑0.45 | 73.94↓1.47 | 46.28↓5.78-54.42↑5.33-47.26↓1.64 | 58.32↓1.90 | 48.28↓0.74 |
| GPT-4o-mini | 76.23↓0.30-67.71↑0.31-67.92↓0.03 | 29.08↓1.00 | 34.24↓0.97 | 33.75↑0.48-45.07↑0.52-36.51↑0.52 | 27.74↓0.41 | 24.64↓0.21 |
| Claude 3 Opus | 65.60↑2.21-85.01↓4.57-72.07↓1.31 | 75.20↓2.19 | 72.09↓2.45 | 51.10↑1.03-74.59↓1.13-58.03↓0.91 | 69.95↓2.41 | 60.72↓1.29 |
| Claude 3.5 Sonnet | 80.36↑2.42-87.93↓0.62-82.33↑0.46 | 74.93↓5.30 | 77.12↓4.45 | 40.74↑5.84-65.18↓0.79-47.75↑3.51 | 58.55↑1.29 | 48.83↑2.60 |
| Claude 3.5 Haiku | 62.69↓8.39-76.01↑0.37-66.02↓5.66 | 40.35↓9.13 | 46.02↓10.1 | 37.42↓6.90-48.72↓4.65-39.91↓6.63 | 29.64↓5.71 | 27.22↓4.48 |
| Gemini 1.5 Pro | 84.64↓4.38-82.25↑4.43-79.73↑1.36 | 76.86↓0.53 | 76.71↑0.49 | 50.40↓4.13-52.33↑0.28-48.42↓2.11 | 69.09↓6.08 | 54.83↓4.45 |
| Llama-3.1-405B | 79.09↑2.72-77.96↓1.60-75.16↑0.36 | 64.25↓0.01 | 67.58↑0.01 | 38.19↓1.08-58.47↑7.38-43.39↑1.12 | 52.98↓2.44 | 43.79↑0.11 |
| Mistral Large | 68.81↑1.40-78.55↓0.18-71.07↑0.89 | 60.82↓2.62 | 63.63↓1.59 | 53.91↓3.05-56.78↓0.32-52.57↓1.94 | 46.24↓1.71 | 46.36↓2.26 |
| DeepSeek-V3 | 86.98↓1.19-73.83↑1.37-75.66↑1.07 | 64.98↑1.84 | 66.50↑1.61 | 47.64↑0.65-42.12↑0.60-42.17↑0.65 | 56.02↑0.77 | 44.09↑0.72 |
| Qwen-Max | 74.59↓5.53-79.86↓1.30-73.79↓2.78 | 60.93↓1.24 | 64.34↓1.59 | 46.49↓0.02-53.59↓0.35-46.87↓0.17 | 45.29↓1.54 | 40.87↓0.23 |
| GLM-4-Plus | 80.28↓1.03-72.81↑0.61-71.17↑0.34 | 69.34↓1.73 | 67.55↓0.43 | 43.50↓1.77-46.93↑0.43-41.64↓0.61 | 55.80↓3.17 | 43.54↓1.26 |
5 Evaluation Setup
Evaluated LLMs.
We evaluate leading LLMs, covering both the most advanced and lightweight variants555Our evaluation began on January, 2025, and all LLMs used the latest official API versions available at that time.: GPT-4o, GPT-4-turbo, GPT-4o-mini, Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku, Gemini 1.5 Pro, Llama-3.1-405B, Mistral Large, DeepSeek-V3, Qwen-Max, GLM-4-Plus. All model sources are listed in Appendix E.
Solution Paradigms.
We benchmark LLMs for KGDS under two interaction paradigms: (1) Single-turn Structured-prompt: A standard prompt with well-structured Li et al. (2023) contains input content, input definition, task description, output definition, and return format (in Appendix D.3). (2) Multi-turn Self-reflection: A second-round self-reflection Shinn et al. (2023) instruction with step-by-step chain-of-thought Wei et al. (2022) reasoning after the structured prompt (in Appendix D.4).
Verifier and Detector.
6 Analysis
In this part, we primarily analyze the commonalities and differences among LLMs in KGDS. Sections §6.1, §6.2, and §6.3 respectively discuss the performance of background summary, opinion summary, and both patterns under structured-prompt. §6.4 explores the impact of LLM self-reflection.
6.1 Background Summary Analysis
LLMs are moderate but imbalanced retrievers for EBS.
From Table 1, we find that most LLMs achieve moderate retrieval performance () and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from and , we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio (Figure 2) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
LLMs prefer expert consistently annotated EBS over their own extractions.
We conduct an investigation to verify whether our EBS is the true gold standard. Specifically, for EBS extracted by LLM that is inconsistent with expert-annotated, we instruct the LLM to perform a second-round feedback judgment to determine which EBS is more aligned with its definition (the instruction is provided in Appendix D.5). As shown in Figure 3, all LLMs unanimously and overwhelmingly prefer our annotations, with the average win rate of , indicating that: (1) Our expert-annotated EBS can be regarded as reliable ground truth. (2) The current LLMs possess latent awareness of optimal retrieval boundaries but struggle to implement them.
LLMs are inadequate generators for ABS.
As shown in Table 1, all LLMs demonstrate unsatisfactory performance (). The weak () reveals that LLMs often omit key facts, while the mediocre () indicates persistent inclusion of irrelevant facts. This dual failure reflects the fundamental deficiencies of LLMs in meeting the requirements of coverage (i.e., completeness) and focus (i.e., conciseness). Moreover, we also observe polarization among LLMs: some prioritize precision at the cost of recall (e.g., Claude 3 Opus), while others attempt to balance both (e.g., GPT-4o). Unlike EBS, we did not find any extreme recall-oriented models, indicating that LLMs tend to be either conservative or balanced in ABS generation.
LLMs are better retrievers than generators for background summary.
As presented in Table 1, all LLMs exhibit higher BSP metrics compared to their KBSAF counterparts777BSP and KBSAF metrics are directly comparable, as paragraph and atomic fact differ only in granularity, and there is no difference in mathematical form between the twin metrics.. This gap between retrieval and generation indicates that LLMs possess stronger in-context recognition capabilities for coarse-grained paragraphs than fine-grained facts.
6.2 Opinion Summary Analysis
Investigating correlation variables influencing AOS quality.
From Table 1, we find that decreases as the background summary quality declines (i.e., ) across all LLMs during pattern transitions. Additionally, as shown in Figures 4 and 5, is highly positively correlated with and among models in independent patterns. These findings indicate that a high-quality background summary facilitates knowledge integration in AOS generation. Meanwhile, individual differences suggest that the model-inherent integration ability is also a key factor affecting AOS quality. For instance, Gemini 1.5 Pro and Claude 3.5 Sonnet show relatively stronger (i.e., above the line in Figure 5) and weaker (i.e., below the line in Figure 4) integration capabilities in ABS-AOS and EBS-AOS patterns, respectively.
As for absolute performance, the most advanced models (highlighted LLMs in Table 1) only achieve an average of across both patterns, indicating that numerous opinions are being incorrectly integrated. Moreover, we observe that lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) exhibit significant flaws in , suggesting that AOS integration is sensitive to parameter scale.
AOS error detection analysis.
As shown in Table 2, a significant proportion of integration errors are concentrated in IRU and IRIC across all LLMs in both patterns, indicating that LLMs struggle to effectively and correctly clarify implicit references during AOS integration. Furthermore, although the ratios of OFI, OSD, and OM errors are relatively lower, their presence still emphasizes the inherent deficiencies of factual inconsistency, sentiment distortion, and incorrect speaker assignment in LLMs.
| Model Name | AOS Error Detection (single-turn structured-prompt and multi-turn self-reflection) | |||||||||
| EBS-AOS Pattern | ABS-AOS Pattern | |||||||||
| OFI | OSD | IRU | IRIC | OM | OFI | OSD | IRU | IRIC | OM | |
| GPT-4o | 8.57↑1.09 | 6.19↓0.88 | 26.67↑2.80 | 55.71↓2.09 | 2.86↓0.92 | 4.07↓1.12 | 2.96↑1.10 | 50.74↓1.66 | 40.37↑1.33 | 1.86↑0.35 |
| GPT-4-turbo | 7.23↓0.38 | 9.24↓1.18 | 40.96↑1.38 | 39.76↓0.65 | 2.81↑0.83 | 6.74↑0.40 | 3.23↑2.33 | 56.33↓0.51 | 31.00↓1.11 | 2.70↓1.11 |
| GPT-4o-mini | 0.79↑0.30 | 0.32↑0.15 | 81.80↓0.08 | 16.93↓0.37 | 0.16↓0.00 | 1.74↓0.79 | 0.79↑0.15 | 73.69↓0.30 | 23.61↑1.11 | 0.17↓0.17 |
| Claude 3 Opus | 6.96↑0.54 | 2.17↓0.09 | 32.17↑0.75 | 56.96↓1.96 | 1.74↑0.76 | 6.61↑0.64 | 1.18↑0.99 | 46.69↑0.05 | 43.19↓2.61 | 2.33↑0.93 |
| Claude 3.5 Sonnet | 6.67↑0.24 | 5.78↓1.78 | 50.22↑2.51 | 35.11↓1.29 | 2.22↑0.32 | 3.00↑3.13 | 3.54↓0.20 | 61.85↓8.65 | 29.97↑5.13 | 1.64↑0.59 |
| Claude 3.5 Haiku | 6.81↓2.01 | 4.09↓0.83 | 55.45↑4.07 | 32.10↓1.23 | 1.55↓0.00 | 3.91↓0.10 | 1.47↓0.41 | 70.52↓2.64 | 23.94↑3.31 | 0.16↓0.16 |
| Gemini 1.5 Pro | 6.28↓0.91 | 7.73↑2.03 | 37.68↑3.78 | 46.86↓3.45 | 1.45↓1.45 | 3.62↑0.05 | 6.52↓0.40 | 63.41↑4.17 | 26.09↓4.38 | 0.36↑0.56 |
| Llama-3.1-405B | 6.92↓1.00 | 1.89↑1.23 | 55.35↑0.10 | 34.59↓0.95 | 1.25↑0.62 | 3.83↑1.18 | 0.47↑1.81 | 69.62↑2.13 | 25.36↓5.77 | 0.72↑0.65 |
| Mistral Large | 6.67↑1.35 | 3.33↓0.39 | 52.78↓0.91 | 36.94↓1.91 | 0.28↑1.86 | 3.73↓0.12 | 2.28↓0.48 | 64.52↑5.82 | 29.05↓5.60 | 0.42↑0.38 |
| DeepSeek-V3 | 4.79↑1.61 | 2.24↓0.22 | 53.35↓5.54 | 38.98↑4.12 | 0.64↑0.03 | 5.99↓0.67 | 2.08↑0.05 | 61.98↓2.94 | 28.91↑3.80 | 1.04↓0.24 |
| Qwen-Max | 7.54↑0.13 | 2.90↑0.79 | 43.48↑1.12 | 45.22↓2.32 | 0.86↑0.28 | 4.47↑0.05 | 3.25↓0.30 | 65.45↑1.94 | 26.22↓1.66 | 0.61↓0.03 |
| GLM-4-Plus | 10.94↓1.09 | 3.02↑1.72 | 40.38↓5.71 | 44.53↑4.74 | 1.13↑0.34 | 5.23↑2.60 | 1.93↑1.99 | 55.65↑0.22 | 35.81↓4.74 | 1.38↓0.07 |
6.3 Overall Performance Analysis
Performance stratification among LLMs.
We observe that metrics exhibit distinct hierarchies in independent patterns and cross-pattern (refer to Figures 6, 8, and 9 in Appendix A). This suggests that the performance of LLMs improves in a stepwise manner rather than continuously as their intelligence advances in KGDS. Moreover, even the best-performing models achieve less than average overall performance across both patterns (see Figure 9), highlighting that KGDS remains a challenging task for current SOTA LLMs.
Cross-pattern stability of LLMs.
By analyzing the performance gap between the two patterns (see Figures 6, 7, and 8 in Appendix A), we find that different LLMs excel in different patterns. For example, Claude 3.5 Sonnet exhibits a significant gap ( ), while Claude 3 Opus demonstrates a relatively smaller gap ( ), indicating stronger cross-pattern stability.
LLMs perform better with EBS-AOS than ABS-AOS for KGDS.
As presented in Table 1, all LLMs achieve superior in EBS-AOS pattern than ABS-AOS, which is attributed to the more complete and accurate background summary and higher-quality opinion summary. Hence, we suggest prioritizing EBS-AOS in real-world KGDS for better implementation.
6.4 Self-Reflection Impact
Self-reflection does not essentially affect the performance of LLMs on KGDS.
From Table 1, we observe that the performance fluctuations (i.e., or ) after self-reflection are limited (the maximum average fluctuation across all metrics is for ). This means that self-reflection does not fundamentally influence the capacities of LLMs in KGDS, revealing two key limitations: (1) LLMs lack sufficient self-evaluation capacities for KGDS. (2) The reflection strategies of LLMs struggle to provide excellent reasoning paths for KGDS.
Self-reflection makes LLMs more conservative or open for background summary.
As shown in Table 1, most LLMs demonstrate polarization in the increases and decreases between and , as well as and . This suggests different reflection strategies: some prioritize precision (e.g., DeepSeek-V3 with , ), while others prioritize recall (e.g., Claude 3 Opus). A few models exhibit simultaneous increases (e.g., GPT-4o with , ) or decreases (e.g., Qwen-Max), indicating more balanced or weaker reflection abilities.
Self-reflection makes LLMs more risk-averse for opinion summary.
As presented in Table 2, most LLMs show a decrease in IRIC and an increase in IRU (e.g., GPT-4o with IRIC , IRU ). This indicates that self-reflection drives LLMs to adopt more cautious strategies when clarifying implicit references — preferring to leave references unclarified rather than risk incorrect clarification. However, excessive conservatism in clarifying implicit references harms AOS quality, as evidenced by the widespread decrease in in Table 1.
7 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than average performance across both patterns. These findings highlight the need for future advancements in LLM capabilities for coarse-grained retrieval, fine-grained generation and knowledge integration.
Limitations
Our study presents several limitations that merit discussion and future exploration:
Domain Generalization.
Knowledge-grounded discussions and their confusing summaries are prevalent in both open-domain and private scenarios. Our benchmark is centered on open-domain news discussions, which may limit its applicability to private scenarios (e.g., internal meetings, medical consultations). However, shared background knowledge and the related discussions in private contexts are challenging to obtain and are less common and representative than our open benchmark. Future work will focus on validating the performance of LLMs in handling KGDS across diverse private scenarios.
Language Diversity.
Our current work is focused on English, creating a monolingual benchmark that may not fully address the challenges of multilingual KGDS. This limitation could impact model performance in languages with different linguistic structures or sociocultural contexts. Nevertheless, English is the most prevalent language in both academic research and real-world LLM applications, making it a reasonable starting point. We plan to explore multilingual and cross-lingual KGDS in future research.
Ethics Statement
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using this benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions. We provide the annotation and evaluation costs as:
Annotation Cost for KGDS BenchMark.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is hours (including news reading and understanding, discussion construction, back-supporting paragraphs, key back-supporting atomic facts, back-nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of $ per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is $ ( hours per sample $ per hour experts per sample samples $).
Evaluation Cost for KGDS BenchMark.
The evaluation cost of API calls is detailed in Table 3.
| Project | Cost ($) |
| Atomic Fact Decomposition | |
| Conflicting Fact Masking | |
| Structured Prompt | |
| Reflection Instruction | |
| Atomic Fact Verification | |
| Atomic Opinion Verification | |
| Fine-Grained Error Detection | |
| Feedback Judgment | |
| Total |
Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We also thank the anonymous reviewers for their constructive feedback that helped improve this work.
Appendix A Overall Performance Visualization
Figures 6 and 8 show the overall performance of EBS-AOS and ABS-AOS. Figures 7 and 9 illustrate the performance gap and average performance.
Appendix B KGDS BenchMark Statistics
News Statistics.
Refer to Table 4.
| Num. | Paras | Tokens |
Discussion Statistics.
Refer to Table 5.
| Num. | Pts. | Uttrs. | Tokens |
Annotation Statistics for EBS.
Among the paragraphs from original news articles, the expert annotation consistency rate is . A total of paragraphs are retained for the final news articles, while are identified as boundary paragraphs and removed. Of the retained, are annotated as back-supporting and as back-nonsupporting.
Annotation Statistics for ABS.
Among the atomic fact units decomposed from the back-supporting paragraphs, are consistently annotated as key back-supporting atomic facts, as non-key back-supporting atomic facts, and as intra-repetitive atomic facts.
Among the atomic fact units from the back-nonsupporting paragraphs, are automatically annotated as back-nonsupporting atomic facts, as masked conflicting atomic facts, and as intra-repetitive atomic facts.
Annotation Statistics for AOS.
A total of clear atomic opinions are expert manually annotated from discussions and their corresponding new articles. Of these, contain clarified implicit references, while do not require clarification. Within the opinions, there are a total of written clarified implicit references.
Appendix C Detailed EBS Evaluation Metrics
Let , , and denote the sets of paragraph indices for LLM-extracted EBS, back-supporting, and back-nonsupporting paragraphs, respectively. The automatic paragraph index matching function is:
where and represent whether the paragraph index is present in .
The BSP Recall, Precision, and F1-score are defined as follows:
Appendix D Prompts and Instructions
D.1 Unified Parameter Settings
In this work, All evaluated LLMs use consistent parameter settings for prompts and instructions: max_token= and temperature=. No other default parameters are modified.
D.2 Atomic Fact Decomposition Instruction
Refer to Figure 11.
D.3 Structured Prompts
D.4 Self-Reflection Instructions
D.5 Feedback Judgment Instruction
Refer to Figure 16.
D.6 Verification Prompts
D.7 Error Detection Instruction
Refer to Figure 18.
Appendix E LLM Sources
OpenAI: GPT-4o, GPT-4-turbo, GPT-4o-mini
https://platform.openai.com/docs/api-reference/introduction
Anthropic: Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku
https://docs.anthropic.com/en/api/getting-started
Google: Gemini 1.5 Pro
https://ai.google.dev/gemini-api/docs
Meta: Llama-3.1-405B
https://www.llmapi.com
Mistral AI: Mistral Large
https://docs.mistral.ai/api/
DeepSeek: DeepSeek-V3
https://api-docs.deepseek.com/zh-cn/
Alibaba: Qwen-Max
https://bailian.console.aliyun.com/
ZhipuAI: GLM-4-Plus
https://www.bigmodel.cn/dev/api/normal-model/glm-4