### v2

# What Are They Talking About? A Benchmark of Knowledge-Grounded Discussion Summarization

Abstract
Traditional dialogue summarization primarily focuses on dialogue content, assuming it comprises adequate information for a clear summary. However, this assumption often fails for discussions grounded in shared background, where participants frequently omit context and use implicit references. This results in summaries that are confusing to readers unfamiliar with the background. To address this, we introduce Knowledge-Grounded Discussion Summarization (KGDS), a novel task that produces a supplementary background summary for context and a clear opinion summary with clarified references. To facilitate research, we construct the first KGDS benchmark, featuring news-discussion pairs and expert-created multi-granularity gold annotations for evaluating sub-summaries. We also propose a novel hierarchical evaluation framework with fine-grained and interpretable metrics. Our extensive evaluation of 12 advanced large language models (LLMs) reveals that KGDS remains a significant challenge. The models frequently miss key facts and retain irrelevant ones in background summarization, and often fail to resolve implicit references in opinion summary integration. [^1] Our benchmark is available at zhouweixiao/KGDS
What Are They Talking About?
A Benchmark of Knowledge-Grounded Discussion Summarization
[…]
1 Introduction
Dialogue summarization aims to distill key topics and interactions from a dialogue into a concise summary Kirstein et al. (2024); Rennard et al. (2023); Jia et al. (2023). Conventional paradigm relies primarily on dialogue content, whether in benchmark construction Gliwa et al. (2019); Chen et al. (2021); Zhu et al. (2021), methodologies Zhou et al. (2023); Tian et al. (2024); Lu et al. (2025); Zhu et al. (2025), or evaluation Wang et al. (2022); Gao and Wan (2022); Zhu et al. (2023); Tang et al. (2023, 2024b); Liu et al. (2024b); Ramprasad et al. (2024). These efforts implicitly assume that "the dialogue itself contains sufficient information to generate a clearly understandable summary for readers."
However, we find this assumption has fundamental limitations and often fails, particularly when participants discuss shared background knowledge they are already familiar with. Such discussions exhibit two main traits: (1) Information Omission and Implicit Reference: Participants naturally skip mutually known details and frequently use pronouns or phrases to refer to entities or facts within the contextual background. (2) Personal Opinion: Unlike simple information interactions in general dialogues, these discussions focus on exchanging viewpoints, with participants expressing personal opinions from various perspectives. These characteristics make understanding the discussion heavily reliant on background knowledge. Consequently, traditional dialogue summarization paradigm inherits this dependency, leading to confusion for outside readers unfamiliar with the context, leaving them wondering: "What are they talking about?". We present an illustrative example in Figure 1.
To address this problem, we introduce KGDS, a novel task designed to combine shared background knowledge with discussion content to create reader-centered summaries. We argue that a successful KGDS summary must achieve two complementary objectives: (1) to bridge readers’ knowledge gaps by providing the necessary background information that supports the discussion; and (2) to present readers with clear participant opinions by clarifying the implicit references within them. Thus, we formalize these requirements by modeling the task output as a Background Summary and an Opinion Summary. The background summary, which retrieves or condenses relevant background information, can be either Extractive or Abstractive, while the opinion summary integrates clarified opinions and is inherently Abstractive. In Figure 1, we provide a visual example of our task paradigms.
Figure 1: An overview example. Gray Blocks in shared background knowledge denote crucial background details omitted by participants during discussion. Purple Blocks in discussion content indicate referential pronouns or phrases. Blue Blocks in discussion summary represent content that may cause confusion for outside readers. Compared to traditional dialogue summarization, KGDS achieves better reader preference by providing a supplementary background summary and a clear opinion summary, in which Cyan Blocks highlight clarified implicit references.
To advance research in this field, we construct the first KGDS benchmark, situated in a common and realistic scenario of news discussions. It consists of event-rich news articles, each paired with a human-authored discussion. To enable robust assessment, we develop multi-granularity gold evaluation components, which are annotated by experts under strict consistency controls. For the extractive background summary, we annotate coarse-grained supporting and nonsupporting paragraphs. For the abstractive background summary, we create two sets of finest-grained atomic facts Min et al. (2023), including key supporting and nonsupporting facts. For the opinion summary, we introduce clear atomic opinions as basic evaluation units that are minimized and have implicit references clarified.
Furthermore, we propose a novel hierarchical framework for evaluation. For each summarization paradigm, we first evaluate its sub-summaries separately and then aggregate their scores to derive a paradigm-level score. At the sub-summary level, our framework assesses multiple dimensions. For the two types of background summaries, we evaluate their coverage, focus, and overall quality by comparing them against annotated paragraphs or atomic facts, using paragraph index matching or LLM-based fact verification Wei et al. (2024). For the opinion summary, we assess overall quality by checking its coverage of annotated atomic opinions and identify fine-grained integration errors by classifying failure types for uncovered opinions. We also conduct human evaluation and show a strong correlation with our automatic metrics.
We evaluate 12 advanced LLMs on our benchmark under the structured-prompt Li et al. (2023) and self-reflection Shinn et al. (2023) settings. Our comprehensive results demonstrate that KGDS remains a substantial challenge, with even the top-performing models achieving an average score below 69%. We also identify several key weaknesses: LLMs struggle with the coverage-focus trade-off in background summary retrieval, frequently miss key facts during background summary generation, and fail to resolve implicit references in opinion summary integration. Furthermore, the limited effectiveness of self-reflection indicates that current LLMs lack sufficient self-correction abilities for this task. These findings highlight key bottlenecks and provide concrete directions for future advancements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
2 Task Formulation
Let $K$ denote the shared background knowledge among participants and $D$ represent the discussion grounded in $K$. The objective of KGDS is to provide a supplementary background summary and a clear opinion summary by integrating $K$ and $D$. We define two summarization paradigms based on the type of background summary.
EBS-AOS Paradigm.
The output comprises an extractive background summary (EBS) and an abstractive opinion summary (AOS):
$B_{e},O\leftarrow f(K,D),\;B_{e}\subseteq K$ (1)
Here, $f$ is the summarizer. $B_{e}$ is the EBS, defined as extractive background supporting chunks for $D$ from $K$. Chunks are text sequences of a predefined granularity (e.g., sentences, paragraphs, etc.). $O$ is the AOS, defined as clear personal opinions of the participants with clarified implicit references.
ABS-AOS Paradigm.
The output contains an abstractive background summary (ABS) and an abstractive opinion summary (AOS):
$B_{a},O\leftarrow f(K,D)$ (2)
The definitions of $f$ and $O$ follow Eq. (1). $B_{a}$, the ABS, is defined as abstractive background supporting information for $D$ from $K$.
3 Benchmark Construction
3.1 Preliminary
Scenario Setting.
We establish our benchmark scenario as a two-participant discussion of news content for two reasons. First, news discussions are highly prevalent in daily life, making them more representative than private scenarios such as internal meetings and medical consultations. Second, news summarization Goyal et al. (2022); Zhang et al. (2024); Liu et al. (2024a) is a well-established subfield of automatic summarization research.
News Collection.
We collect 100 multi-domain (i.e., business, sports, and world) event-rich news articles from Google News [^2] https://news.google.com as shared background knowledge, with a time cutoff of Oct. 2024. We preserve the original news paragraph structure and define paragraph-as-chunk as the minimum extraction granularity under the EBS-AOS paradigm.
Expert Annotators.
To ensure high-quality data, we recruit four PhD candidates specializing in NLP for our annotation tasks. Each pair of experts forms a collaborative group to conduct the full-process annotation, which includes discussion generation and the creation of multi-granularity evaluation components for background and opinion summaries.
3.2 Annotation
Human Discussion.
This construction follows the sequence of read, understand, then discuss. For each news article, we require two experts to independently read and thoroughly comprehend its content. This preparatory step aims to ensure background consistency between the participants. Afterward, they engage in a discussion to exchange viewpoints. The entire process is open-ended, meaning the discussion initiator is chosen randomly, and the discussion topics can encompass any events, facts, or detailed information within the news article.
Data Scale Avg. Tokens
KGDS Task Inputs
News-Discussion Pair 100 729.5
EBS Evaluation Components
Supporting Paragraph 432 43.8
Nonsupporting Paragraph 1,005 42.6
ABS Evaluation Components
Key Supporting Atomic Fact 1,638 11.3
Nonsupporting Atomic Fact 4,996 11.4
AOS Evaluation Components
Clear Atomic Opinion 873 19.4
Clarified Implicit References 1,113 4.1
Table 1: Benchmark statistics. For three types of sub-summaries, we annotate evaluation components at multiple granularities, including coarse-grained paragraphs, fine-grained facts and opinions. The green boxes highlight the average lengths of fine-grained annotations.
Paragraph-Level Annotation for EBS.   An ideal EBS should include all paragraphs that support the discussion and exclude any that do not. To create gold labels for this, two experts independently classify each paragraph from the source news as either supporting or nonsupporting. We perform a consistency control, retaining only paragraphs where both experts agree on the label. Paragraphs with conflicting annotations are considered ambiguous and are removed from the news source to ensure data clarity. Our statistics show that out of a total of 1,696 paragraphs, 1,437 (84.7%) are annotated consistently. Table 1 provides more statistics.
Atomic-Fact-Level Annotation for ABS.  Unlike coarse-grained paragraph annotation, a model-generated ABS is condensed and requires fine-grained, key-point-focused labeling. Inspired by the minimal granularity of atomic facts Liu et al. (2023b); Tang et al. (2024a), we argue that an ideal ABS should cover all key supporting facts while filtering out nonsupporting ones. Motivated by this, we create these two distinct atomic fact sets.
First, we utilize GPT-4o [^3] Version: gpt-4o-2024-08-06 to decompose supporting paragraphs into candidate atomic facts based on principles of indivisibility, independence, and declarativity (see Appendix F.2). Two experts then independently classify each fact as either key or non-key. Through a consistency check, only facts unanimously labeled as key by both experts are considered as the gold standard, resulting in 1,638 key supporting atomic facts.
Person1 thinks that it was wise of Rees-Zammit (him)
to sign with the Kansas City Chiefs (this team).
Person2 thinks that worrying about Jersey’s (their)
economic diversification is an overconcern.
Table 2: Examples of clear atomic opinions. Each opinion is an indivisible fine-grained basic unit with clarified implicit references. The content in parentheses indicates the original referential pronouns and phrases.
Second, we use the same decomposition method to nonsupporting paragraphs to generate an initial set of nonsupporting facts. From this set, we then mask [^4] The masking process is consistent with automatic fact verification Tang et al. (2024a), and the masked conflicting facts are not considered in ABS evaluation. any conflicting facts, defined as those that could also be inferred from supporting paragraphs. Such conflicts typically arise when identical or similar facts appear in both supporting and nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking these facts, we ensure that the key supporting and nonsupporting fact sets remain non-overlapping, thereby avoiding external verification issues during evaluation. This process yields 4,996 nonsupporting atomic facts, with conflicting facts proving to be sparse at just 176 instances (3.39%).
Atomic-Opinion-Level Annotation for AOS.  According to the task setup, an effective AOS must clearly present participant opinions by clarifying implicit references. To enable fine-grained evaluation, we introduce the clear atomic opinion as the basic unit, which is both minimal and has implicit references clarified. Table 2 provides two examples. We argue that the quality of an AOS can be assessed by measuring its coverage of these atomic opinions. For annotation, we first require experts to extract the main opinions from their respective utterances in the discussion. They then identify referential pronouns and phrases within these opinions and clarify them through anaphora resolution or information supplementation to produce clear opinions. Finally, the experts decompose these clear opinions into atomic opinion units, following the principles of indivisibility and independence. This annotation process is fully expert-authored. We create 873 clear atomic opinions, among which 800 contain clarified implicit references, while 73 require no clarification. Table 1 provides statistics.
Figure 2: An overview example of our evaluation framework. It comprehensively and accurately evaluates sub-summaries and summarization paradigms through fine-grained, interpretable metrics and hierarchical aggregation.
4 Evaluation Framework
Our framework is comprehensive and hierarchical. For each of the two summarization paradigms (§2), we independently evaluate the sub-summaries and then aggregate their quality to assess overall performance at the paradigm level. The evaluation methods are interpretable, and the metrics are fine-grained. Figure 2 shows an example. Below, we first describe the evaluation dimensions [^5] We do not focus on fluency and coherence because current LLMs perform well on these dimensions Song et al. (2025). at the sub-summary and paradigm levels (§4.1), and then introduce our automatic methods and metrics (§4.2).
4.1 Dimensions
Background Summary.
For the two variants of background summaries, we measure the following dimensions of system responses:
  • Coverage: Whether the model-extracted or -generated summary fully covers the supporting information (i.e., paragraphs or key facts).
  • Focus: Whether the summary focuses on supporting information while excluding nonsupporting information.
  • Overall Quality: A holistic assessment of the summary, considering both coverage and focus.
Opinion Summary.
We evaluate the output performance according to the following dimensions [^6] Unlike background summaries, opinion summaries are inherently integrative. The space of opinions that could be wrongly integrated is effectively unbounded, making it infeasible to construct an exhaustive set of all incorrect atomic opinions to be excluded. Therefore, measuring a Focus dimension is impractical. Instead, we identify integration errors in the summary for uncovered clear atomic opinions.:
  • Overall Quality: Whether the summary successfully integrates clarified opinions, enabling it to cover all clear atomic opinions.
  • Integration Error: For uncovered clear atomic opinions, identify the errors in the summary.
Summarization Paradigm.
We assess the overall quality of each summarization paradigm:
  • Overall Quality: How well the background and opinion summaries work together.
4.2 Methods and Metrics
EBS Evaluation.  We evaluate its quality by matching the paragraph indices from the system output against our annotations of supporting and nonsupporting paragraphs. We utilize Supporting Paragraph Recall, Precision, and F1-score to measure the coverage, focus, and overall quality of the summary, respectively.
ABS Evaluation.  We employ an LLM-based verifier Wei et al. (2024) to check if the summary entails our annotated atomic facts. This process determines whether the summary includes key supporting facts while excluding nonsupporting ones. We use three fine-grained metrics, Key Supporting Atomic Fact Recall, Precision, and F1-score, to quantify the three dimensions of the summary. We provide detailed formulas in Appendix C.1.
AOS Evaluation.  Similar to ABS, we utilize an opinion verifier to measure its coverage of our annotated atomic opinions. We utilize Clear Atomic Opinion Recall to assess its overall quality. For each uncovered atomic opinion, we perform LLM-based error detection to identify the specific integration failure. Specifically, we define five fine-grained error types: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion. Detailed formulas and error definitions are provided in Appendix C.2 and F.6.
Paradigm-Level Quality Evaluation.  A successful KGDS output requires high quality in both its background and opinion summaries. To capture this dependency, we evaluate the paradigm-level quality using the Geometric Mean of the two sub-summary overall scores. The multiplicative nature of this metric ensures a high score is only achieved when both components are strong, reflecting their joint fulfillment of the task goal. Moreover, root normalization maintains dimensional consistency between the paradigm-level score and its parts.
Model Name KGDS BenchMark (single-turn structured-prompt and multi-turn self-reflection)
EBS-AOS Paradigm ABS-AOS Paradigm
SP (R-P-F1) CAO (R) PLQ (GM) KSAF (R-P-F1) CAO (R) PLQ (GM)
GPT-4o 73.39↑1.56 88.10↑0.34 78.12↑1.27 76.18↑0.54 76.24↑1.04 63.57↓1.08 61.73↑1.62 58.34↑0.60 69.42↓0.21 61.09↑0.47
GPT-4-turbo 73.03↓7.72 87.42↑1.67 77.14↓3.80 72.73↑0.45 73.94↓1.47 46.28↓5.78 54.42↑5.33 47.26↓1.64 58.32↓1.90 48.28↓0.74
GPT-4o-mini 76.23↓0.30 67.71↑0.31 67.92↓0.03 29.08↓1.00 34.24↓0.97 33.75↑0.48 45.07↑0.52 36.51↑0.52 27.74↓0.41 24.64↓0.21
Claude 3 Opus 65.60↑2.21 85.01↓4.57 72.07↓1.31 75.20↓2.19 72.09↓2.45 51.10↑1.03 74.59↓1.13 58.03↓0.91 69.95↓2.41 60.72↓1.29
Claude 3.5 Sonnet 80.36↑2.42 87.93↓0.62 82.33↑0.46 74.93↓5.30 77.12↓4.45 40.74↑5.84 65.18↓0.79 47.75↑3.51 58.55↑1.29 48.83↑2.60
Claude 3.5 Haiku 62.69↓8.39 76.01↑0.37 66.02↓5.66 40.35↓9.13 46.02↓10.1 37.42↓6.90 48.72↓4.65 39.91↓6.63 29.64↓5.71 27.22↓4.48
Gemini 1.5 Pro 84.64↓4.38 82.25↑4.43 79.73↑1.36 76.86↓0.53 76.71↑0.49 50.40↓4.13 52.33↑0.28 48.42↓2.11 69.09↓6.08 54.83↓4.45
Llama-3.1-405B 79.09↑2.72 77.96↓1.60 75.16↑0.36 64.25↓0.01 67.58↑0.01 38.19↓1.08 58.47↑7.38 43.39↑1.12 52.98↓2.44 43.79↑0.11
Mistral Large 68.81↑1.40 78.55↓0.18 71.07↑0.89 60.82↓2.62 63.63↓1.59 53.91↓3.05 56.78↓0.32 52.57↓1.94 46.24↓1.71 46.36↓2.26
DeepSeek-V3 86.98↓1.19 73.83↑1.37 75.66↑1.07 64.98↑1.84 66.50↑1.61 47.64↑0.65 42.12↑0.60 42.17↑0.65 56.02↑0.77 44.09↑0.72
Qwen-Max 74.59↓5.53 79.86↓1.30 73.79↓2.78 60.93↓1.24 64.34↓1.59 46.49↓0.02 53.59↓0.35 46.87↓0.17 45.29↓1.54 40.87↓0.23
GLM-4-Plus 80.28↓1.03 72.81↑0.61 71.17↑0.34 69.34↓1.73 67.55↓0.43 43.50↓1.77 46.93↑0.43 41.64↓0.61 55.80↓3.17 43.54↓1.26
Table 3: Main evaluation results. SP, KSAF, CAO, and PLQ represent Supporting Paragraph, Key Supporting Atomic Fact, Clear Atomic Opinion, and Paradigm-Level Quality, respectively. R, P, F1, and GM denote Recall, Precision, F1-score, and Geometric Mean, respectively. All reported metrics are macro-averaged (
). $\uparrow$ and $\downarrow$ indicate performance increases and decreases, respectively, after self-reflection following the structured-prompt. For each overall metric (i.e., $\mathrm{SP}_{\mathrm{F_{1}}}$, $\mathrm{KSAF}_{\mathrm{F_{1}}}$, $\mathrm{CAO}_{\mathrm{R}}$, and $\mathrm{PLQ}_{\mathrm{GM}}$), we highlight the top-3 performing models.
5 Experimental Setup
System Selection.
We select 12 LLMs as summarizers, covering the most advanced and lightweight variants [^7] Our evaluation began on January, 2025, and all LLMs used the latest official API versions available at that time.: GPT-4o, GPT-4-turbo, GPT-4o-mini, Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku, Gemini 1.5 Pro, Llama-3.1-405B, Mistral Large, DeepSeek-V3, Qwen-Max, GLM-4-Plus. All model sources are listed in Appendix E.
Prompt Engineering.
We benchmark LLMs for KGDS under two interaction patterns: (1) Single-turn Structured-prompt: A well-structured standard prompt Li et al. (2023) that includes input content, input definition, task description, output definition, and return format (Appendix F.3). (2) Multi-turn Self-reflection: A second-round self-reflection instruction Shinn et al. (2023) with step-by-step chain-of-thought reasoning Wei et al. (2022) following the structured prompt (Appendix F.4).
Verifier and Detector.
We use GPT-4o [^8] Version: gpt-4o-2024-11-20 to perform fact and opinion verification as well as error detection due to its excellent consistency with human judgment Song et al. (2024a, b). All prompts are provided in Appendices F.5 and F.6.
Figure 3: Paragraph retrieval ratios (
[…]
6 Analysis
In this section, we primarily reveal the challenges LLMs face in KGDS and analyze the commonalities and differences among them. Sections §6.1, §6.2, and §6.3 respectively discuss the performance in background summary, opinion summary, and both paradigms under the structured-prompt setting. §6.4 explores the impact of LLM self-reflection.
6.1 Background Summary
LLMs are moderate but imbalanced retrievers for EBS.
From Table 3, we find that most LLMs achieve moderate retrieval performance ($\mathrm{SP}_{\mathrm{F_{1}}}\in[71,82]$) and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from $\mathrm{SP}_{\mathrm{R}}$ and $\mathrm{SP}_{\mathrm{P}}$, we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio [^9] The ratio of the number of retrieved paragraphs to the total number of paragraphs. (Figure 3) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
LLMs are inadequate generators for ABS.
As presented in Table 3, all LLMs exhibit unsatisfactory performance ($\mathrm{KSAF}_{\mathrm{F_{1}}}\in[37,58]$). The low $\mathrm{KSAF}_{\mathrm{R}}$ (average of 46.08%) reveals that LLMs often omit key facts, while the weak $\mathrm{KSAF}_{\mathrm{P}}$ (average of 54.99%) indicates persistent inclusion of irrelevant facts. This dual-failure reflects the fundamental deficiencies of LLMs in meeting the requirements of coverage and focus. We also observe polarization among LLMs: some prioritize precision at the cost of recall (e.g., Claude 3.5 Sonnet), while others attempt to balance both (e.g., GPT-4o). Unlike EBS, we do not find any extreme recall-oriented models, indicating that LLMs tend to be either conservative or balanced in ABS generation.
6.2 Opinion Summary
Investigating correlation variables influencing AOS quality.
From Table 3, we find that $\mathrm{CAO}_{\mathrm{R}}$ decreases as the background summary quality declines (i.e., $\mathrm{SP}_{\mathrm{F_{1}}}\rightarrow\mathrm{KSAF}_{\mathrm{F_{1}}}$) across all LLMs when switching paradigms. Furthermore, as shown in Figures 4 and 5, $\mathrm{CAO}_{\mathrm{R}}$ is highly positively correlated with $\mathrm{SP}_{\mathrm{F_{1}}}$ and $\mathrm{KSAF}_{\mathrm{F_{1}}}$ among LLMs in both independent paradigms. These findings indicate that a high-quality background summary facilitates opinion integration. Meanwhile, individual differences suggest that the model-inherent integration ability is also a key factor affecting AOS performance. For instance, Claude 3.5 Sonnet and Gemini 1.5 Pro show relatively weaker (i.e., below the line in Figure 4) and stronger (i.e., above the line in Figure 5) integration capabilities in EBS-AOS and ABS-AOS paradigms, respectively.
Figure 4: Visualization and metrics of the correlation between $\mathrm{SP}_{\mathrm{F_{1}}}$ and $\mathrm{CAO}_{\mathrm{R}}$ under the EBS-AOS paradigm.
Nevertheless, even the most advanced models achieve only an average of 72% $\mathrm{CAO}_{\mathrm{R}}$ across both paradigms, indicating that numerous opinions are being incorrectly integrated. Moreover, lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) exhibit significant performance flaws, suggesting that opinion integration is sensitive to model scale.
AOS error distribution analysis.
As illustrated in Table 5, a significant ratio of errors is concentrated in IRU and IRIC across all LLMs under both paradigms, indicating that LLMs struggle to clarify implicit references effectively and correctly during opinion integration. Moreover, although the ratios of OFI, OSD, and OM errors are relatively lower, their presence still highlights inherent deficiencies such as factual inconsistency, sentiment distortion, and incorrect participant assignment in LLMs.
6.3 Summarization Paradigm
Performance stratification among LLMs.
We observe that the overall metric $\mathrm{PLQ}_{\mathrm{GM}}$ exhibits distinct hierarchies in independent paradigms and cross-paradigm (Figures 6, 8, and 9 in Appendix A). This suggests that the performance of LLMs improves in a stepwise manner rather than continuously as their intelligence advances in KGDS. Nevertheless, even the best-performing models achieve less than 69% average performance across both paradigms (see Figure 9), highlighting that KGDS remains a challenging task for current LLMs.
Our evaluation framework aligns well with human judgment.
To validate this, we conduct a human evaluation of overall paradigm-level quality [^10] We focus on paradigm-level validation because the KGDS task’s success depends on the complementarity of both sub-summaries. Furthermore, the individual sub-summary evaluations are already grounded in expert annotations and interpretable, fine-grained verification methods that are known to correlate well with human assessment Song et al. (2024a).. We randomly sample 120 summary pairs (10 per model and 5 per paradigm) from all outputs, and two human evaluators rate their overall quality on a 1-5 Likert scale Joshi et al. (2015), using the average as the final human score. For comparison, we select three LLMs (i.e., GPT-4o, Llama-3.1-405B, DeepSeek-V3) as LLM-as-a-judge baselines to directly score the summaries, with prompt settings following G-Eval Liu et al. (2023a). Table 4 presents the Pearson, Spearman, and Kendall correlation coefficients. Our framework achieves significantly higher correlation with human judgments than the baselines, demonstrating the effectiveness of our fine-grained, hierarchical approach.
Figure 5: Vis. and metrics of the correlation between $\mathrm{KSAF}_{\mathrm{F_{1}}}$ and $\mathrm{CAO}_{\mathrm{R}}$ under the ABS-AOS paradigm.
Metric Pearson Spearman Kendall
LLM-as-a-judge
GPT-4o .5273 .5162 .4193
Llama-3.1-405B .3506 .3760 .2881
DeepSeek-V3 .4282 .4417 .3479
Our Framework .6717 .6602 .5829
Table 4: Correlation coefficients between different evaluation methods and human judgments, where our framework demonstrates the best alignment ($p<0.01$).
Model Name AOS Error Detection (single-turn structured-prompt and multi-turn self-reflection)
EBS-AOS Paradigm ABS-AOS Paradigm
OFI OSD IRU IRIC OM OFI OSD IRU IRIC OM
GPT-4o 8.57↑1.09 6.19↓0.88 26.67↑2.80 55.71↓2.09 2.86↓0.92 4.07↓1.12 2.96↑1.10 50.74↓1.66 40.37↑1.33 1.86↑0.35
GPT-4-turbo 7.23↓0.38 9.24↓1.18 40.96↑1.38 39.76↓0.65 2.81↑0.83 6.74↑0.40 3.23↑2.33 56.33↓0.51 31.00↓1.11 2.70↓1.11
GPT-4o-mini 0.79↑0.30 0.32↑0.15 81.80↓0.08 16.93↓0.37 0.16↓0.00 1.74↓0.79 0.79↑0.15 73.69↓0.30 23.61↑1.11 0.17↓0.17
Claude 3 Opus 6.96↑0.54 2.17↓0.09 32.17↑0.75 56.96↓1.96 1.74↑0.76 6.61↑0.64 1.18↑0.99 46.69↑0.05 43.19↓2.61 2.33↑0.93
Claude 3.5 Sonnet 6.67↑0.24 5.78↓1.78 50.22↑2.51 35.11↓1.29 2.22↑0.32 3.00↑3.13 3.54↓0.20 61.85↓8.65 29.97↑5.13 1.64↑0.59
Claude 3.5 Haiku 6.81↓2.01 4.09↓0.83 55.45↑4.07 32.10↓1.23 1.55↓0.00 3.91↓0.10 1.47↓0.41 70.52↓2.64 23.94↑3.31 0.16↓0.16
Gemini 1.5 Pro 6.28↓0.91 7.73↑2.03 37.68↑3.78 46.86↓3.45 1.45↓1.45 3.62↑0.05 6.52↓0.40 63.41↑4.17 26.09↓4.38 0.36↑0.56
Llama-3.1-405B 6.92↓1.00 1.89↑1.23 55.35↑0.10 34.59↓0.95 1.25↑0.62 3.83↑1.18 0.47↑1.81 69.62↑2.13 25.36↓5.77 0.72↑0.65
Mistral Large 6.67↑1.35 3.33↓0.39 52.78↓0.91 36.94↓1.91 0.28↑1.86 3.73↓0.12 2.28↓0.48 64.52↑5.82 29.05↓5.60 0.42↑0.38
DeepSeek-V3 4.79↑1.61 2.24↓0.22 53.35↓5.54 38.98↑4.12 0.64↑0.03 5.99↓0.67 2.08↑0.05 61.98↓2.94 28.91↑3.80 1.04↓0.24
Qwen-Max 7.54↑0.13 2.90↑0.79 43.48↑1.12 45.22↓2.32 0.86↑0.28 4.47↑0.05 3.25↓0.30 65.45↑1.94 26.22↓1.66 0.61↓0.03
GLM-4-Plus 10.94↓1.09 3.02↑1.72 40.38↓5.71 44.53↑4.74 1.13↑0.34 5.23↑2.60 1.93↑1.99 55.65↑0.22 35.81↓4.74 1.38↓0.07
Table 5: Fine-grained error detection results. OFI, OSD, IRU, IRIC, and OM represent the five error types: Opinion Fact Inconsistency, Opinion Sentiment Distortion, Implicit Reference Unclarified, Implicit Reference Incorrectly Clarified, and Opinion Misattribution. All reported metrics are error proportions (
) when clear atomic opinions are not covered by AOS. $\downarrow$ and $\uparrow$ respectively indicate decreases and increases in error ratios after self-reflection.
6.4 Self-Reflection Impact
Self-reflection does not essentially affect the performance of LLMs on KGDS.
[…]
Self-reflection makes LLMs more risk-averse for opinion summary.
[…]
7 Related Work
Dialogue Summarization.  This task spans diverse scenarios, including online chats Gliwa et al. (2019), meetings Hu et al. (2023), and customer service Lin et al. (2021). Existing methods, whether based on feature modeling Chen and Yang (2021); Fang et al. (2022); Lin et al. (2022), pre-training Zou et al. (2021); Zhong et al. (2022), or LLMs Wang et al. (2023a); Zhu et al. (2025), focus heavily on single-source dialogue inputs while neglecting shared background knowledge among participants. Moreover, current evaluation systems Ramprasad et al. (2024); Tang et al. (2024b) also assume single-source inputs. In contrast, our KGDS focuses on multi-source inputs and complementary outputs.
Knowledge-Intensive Dialogue Response.  This task aims to enhance the dialogue process by incorporating external knowledge sources Lewis et al. (2020); Chen et al. (2024). It differs from our KGDS in two key aspects. First, knowledge-intensive dialogue focuses on improving the conversational experience for participants Wang et al. (2023b, 2024), whereas KGDS is designed to generate clear summaries for external readers. Second, the former addresses knowledge gaps between participants, making the dialogue inherently knowledge-intensive Zhang et al. (2020). In contrast, KGDS centers on discussions among participants with shared knowledge, resulting in background-sparse conversations that rely heavily on implicit context.
Summarization Evaluation.  Earlier similarity-based metrics like Rouge Lin (2004), BERTScore Zhang et al. (2019), and MoverScore Zhao et al. (2019) are often misaligned with human judgment. Recently, LLM-based evaluators Chen et al. (2023); Shen et al. (2023); Fu et al. (2024) have been used, yet they still lack interpretability. To address this, some studies Song et al. (2024a); Lee et al. (2024); Yang et al. (2024); Scirè et al. (2024) utilize atomic facts Min et al. (2023) and LLM-based claim verification Song et al. (2024b), achieving more fine-grained evaluation. Among these, FineSurE Song et al. (2024a) is most relevant to our work. However, its metrics lack consistency in evaluation granularity. In contrast, our coverage and focus metrics are both measured at the atomic fact level, ensuring consistent granularity across dimensions.
8 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of 12 leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than 69% average performance across both paradigms.
Limitations
Our current work is focused on English, creating a monolingual benchmark that may not fully address the challenges of multilingual KGDS. This limitation could impact model performance in languages with different linguistic structures or sociocultural contexts. Nevertheless, English is the most prevalent language in both academic research and real-world LLM applications, making it a reasonable starting point. We plan to explore multilingual and cross-lingual KGDS in future research.
Ethics Statement
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using our KGDS benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions. We provide the annotation and evaluation costs as:
Project Cost (\$)
Atomic Fact Decomposition 212121
Conflicting Fact Masking 161616
Structured Prompt 747474
Reflection Instruction 126126126
Atomic Fact Verification 173173173
Atomic Opinion Verification 158158158
Fine-Grained Error Detection 898989
Total 657657657
Table 6: API call costs for the KGDS benchmark.
Annotation Cost for KGDS BenchMark.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is 2.62.62.6 hours (including news reading and understanding, discussion construction, supporting paragraphs, key supporting atomic facts, nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of \$999 per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is \$468046804680 (2.62.62.6 hours per sample $\times$ \$999 per hour $\times$ 222 experts per sample $\times$ 100100100 samples $=$ \$468046804680). The evaluation cost of API calls is detailed in Table 6.
Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We gratefully acknowledge the research funding provided by Fanyu AI Laboratory. We thank the anonymous reviewers for their constructive feedback that helped improve this work.
Appendix A Paradigm-Level Performance Visualization
Figure 6 and Figure 8 illustrate the overall performance of each model under the EBS-AOS and ABS-AOS paradigms, respectively. Figure 9 presents the average performance of each model across both paradigms. Figure 7 reveals the cross-paradigm stability of each model by quantifying the performance gap between the two paradigms.
Figure 6: Overall performance under the EBS-AOS paradigm. Models are stratified into three tiers: Tier-1 ([72,77][72,77][ 72 , 77 ]), Tier-2 ([64,68][64,68][ 64 , 68 ]), and Tier-3 ([34,46][34,46][ 34 , 46 ]).
Figure 7: Overall performance gap between the two paradigms. A larger gap indicates lower stability.
Appendix B Further Analysis
LLMs are better retrievers than generators for background summary.
As presented in Table 3, all LLMs exhibit higher $\mathrm{SP}$ metrics compared to their $\mathrm{KSAF}$ counterparts. This gap between retrieval and generation indicates that LLMs possess stronger in-context recognition capabilities for coarse-grained paragraphs than fine-grained facts.
Cross-paradigm stability of LLMs.
[…]
LLMs perform better with EBS-AOS than ABS-AOS for KGDS.
As presented in Table 3, all LLMs achieve superior $\mathrm{PLQ}_{\mathrm{GM}}$ in the EBS-AOS paradigm compared to ABS-AOS, due to the more complete and accurate background summaries and higher-quality opinion summaries. Therefore, we suggest prioritizing EBS-AOS in real-world KGDS for more effective implementation.
Self-reflection makes LLMs more conservative or open for background summary.
[…]
Appendix C Detailed Formulas of Evaluation Metrics
C.1 ABS Evaluation Metrics
Let $\mathcal{B}$ denote the LLM-generated ABS, and let $\mathcal{K}=\{k_{i}\}_{i=1}^{m}$ and $\mathcal{N}=\{n_{j}\}_{j=1}^{p}$ represent the sets of $m$ key supporting and $p$ nonsupporting atomic facts, respectively. We define the fact verification function $\phi$ as:
$\phi:(\mathcal{B},f)\rightarrow\{0,1\},\;f\in\mathcal{K}\cup\mathcal{N}$ (3)
where $\phi(\mathcal{B},f)=1$ if the fact $f$ can be inferred from $\mathcal{B}$, and 0 otherwise.
Figure 8: Overall performance under the ABS-AOS paradigm. Models are stratified into three tiers: Tier-1 ([55,61][55,61][ 55 , 61 ]), Tier-2 ([41,49][41,49][ 41 , 49 ]), and Tier-3 ([25,27][25,27][ 25 , 27 ]).
Figure 9: Average overall performance across both paradigms. Models are stratified into three tiers: Tier-1 ([61,69][61,69][ 61 , 69 ]), Tier-2 ([53,56][53,56][ 53 , 56 ]), and Tier-3 ([29,37][29,37][ 29 , 37 ]).
KSAF Recall.
This metric quantifies the coverage of key supporting atomic facts in the ABS:
$\mathrm{KSAF}_{\mathrm{R}}=\frac{1}{m}\sum_{i=1}^{m}\phi(\mathcal{B},k_{i})$ (4)
KSAF Precision.
This metric measures the focus of the ABS on key supporting atomic facts:
$\mathrm{KSAF}_{\mathrm{P}}=\frac{\sum_{k\in\mathcal{K}}\phi(\mathcal{B},k)}{\sum_{f\in\mathcal{K}\cup\mathcal{N}}\phi(\mathcal{B},f)}$ (5)
KSAF F1-score.
This metric evaluates the overall quality of the ABS by comprehensively considering both coverage and focus:
$\mathrm{KSAF}_{\mathrm{F_{1}}}=2\cdot\frac{\mathrm{Precision}\cdot\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}$ (6)
C.2 AOS Evaluation Metrics
Let $\mathcal{O}$ denote the LLM-generated AOS, and let $\mathcal{C}=\{c_{i}\}_{i=1}^{n}$ represent the set of $n$ clear atomic opinions. We define the opinion verification function $\psi$ as:
$\psi:(\mathcal{O},c)\rightarrow\{0,1\},\;c\in\mathcal{C}$ (7)
where $\psi(\mathcal{O},c)=1$ if the opinion $c$ can be inferred from $\mathcal{O}$, and 0 otherwise.
CAO Recall.
This metric quantifies the coverage of clear atomic opinions in the AOS:
$\mathrm{CAO}_{\mathrm{R}}=\frac{1}{n}\sum_{i=1}^{n}\psi(\mathcal{O},c_{i})$ (8)
Appendix D Detailed Benchmark Statistics
News Statistics.
Refer to Table 7.
[…]
100100100 14.414.414.4 617.5617.5617.5
[…]
Discussion Statistics.
Refer to Table 8.
[…]
100100100 2.02.02.0 4.14.14.1 112.0112.0112.0
[…]
Annotation Statistics for EBS.
[…]
Annotation Statistics for ABS.
Among the 242824282428 atomic fact units decomposed from the 432432432 supporting paragraphs, 163816381638 are consistently annotated as key supporting atomic facts, 780780780 as non-key atomic facts, and 101010 as intra-repetitive atomic facts.
Among the 518751875187 atomic fact units from the 100510051005 nonsupporting paragraphs, 499649964996 are automatically annotated as nonsupporting atomic facts, 176176176 as masked conflicting facts, and 151515 as intra-repetitive atomic facts.
Annotation Statistics for AOS.
A total of 873873873 clear atomic opinions are expert manually annotated. Of these, 800800800 contain clarified implicit references, while 737373 do not require clarification. Within the 800800800 opinions, there are a total of 111311131113 written clarified implicit references.
Appendix E LLM Sources
OpenAI [^11] https://platform.openai.com/docs/api-reference/introduction: GPT-4o, GPT-4-turbo, GPT-4o-mini
Anthropic [^12] https://docs.anthropic.com/en/api/getting-started: Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku
Google [^13] https://ai.google.dev/gemini-api/docs: Gemini 1.5 Pro
Meta [^14] https://www.llmapi.com: Llama-3.1-405B
Mistral AI [^15] https://docs.mistral.ai/api/: Mistral Large
DeepSeek [^16] https://api-docs.deepseek.com/zh-cn/: DeepSeek-V3
Alibaba [^17] https://bailian.console.aliyun.com/: Qwen-Max
Appendix F Prompts and Instructions
F.1 Unified Parameter Settings
In this work, all 12 evaluated LLMs use consistent parameter settings for prompts and instructions: max_token=4096 and temperature=0. No other default parameters are modified.
F.2 Atomic Fact Decomposition Instruction
This instruction is used during the benchmark construction phase to guide a large language model in decomposing news paragraphs into fine-grained, atomic fact units. The instruction specifies three core principles for decomposition: indivisibility, independence, and declarativity. This ensures that the resulting fact units are minimal in granularity and informationally complete (see Figure 10).
F.3 Structured Prompts
These prompts serves as the core prompt for evaluating the performance of different LLMs on the KGDS task. It provides the models with standard inputs, including the shared background knowledge and the knowledge-grounded discussion, and clearly defines the objectives and requirements for the two sub-summaries under both summarization paradigms. See Figures 11 and 12 for EBS-AOS and ABS-AOS paradigms, respectively.
F.4 Self-Reflection Instructions
These instructions are used in the multi-turn self-reflection evaluation setting. After the model generates its initial summary, these instructions guide it to critically review its output, checking whether the generated background and opinion summaries strictly adhere to the task definitions. The instructions require the model to provide a detailed chain of thought in a step-by-step manner and, based on this reflection, generate a refined summary. This process is designed to explore and evaluate the model’s self-correction capabilities on the KGDS task. See Figures 13 and 14 for the EBS-AOS and ABS-AOS paradigms, respectively.
F.5 Verification Prompts
These prompts are a core component of our evaluation framework, used to automatically verify the quality of the generated summaries. We utilize a powerful verifier LLM (GPT-4o in our work) to execute these prompts, determining whether the model-generated summary contains the key information from the gold standard. Specifically, the fact verification prompt (see Figure 15) is used to assess the recall and precision of the abstractive background summary against key supporting atomic facts, while the opinion verification prompt (see Figure 16) evaluates the coverage of the abstractive opinion summary against clear atomic opinions.
F.6 Error Detection Instruction
This instruction is used for the fine-grained error analysis of the abstractive opinion summary. When the verifier determines that the summary has failed to cover a gold-standard clear atomic opinion, this instruction is invoked to attribute the failure to one of five predefined, specific error types (e.g., opinion misattribution, implicit reference incorrectly clarified). This approach allows us to gain deep insights into the specific weaknesses of each model in integrating opinions and clarifying references. We provide the instruction in Figure 17.
Figure 10: Atomic fact decomposition instruction.
Figure 11: Structured prompt for KGDS EBS-AOS paradigm.
Figure 12: Structured prompt for KGDS ABS-AOS paradigm.
Figure 13: Self-reflection instruction for KGDS EBS-AOS paradigm.
Figure 14: Self-reflection instruction for KGDS ABS-AOS paradigm.
Figure 15: Atomic fact verification prompt. The number of facts ($n$) dynamically changes at the paragraph-level.
Figure 16: Atomic opinion verification prompt.
Figure 17: AOS error detection instruction.